OpenStudy (anonymous):

When tossing 4 coins simultaneously,find the probability at least 1 head is showing?

OpenStudy (andriod09):

there are 4 coins. Each coin has a \[\frac{1} { 2 }\]chance. you got it so far?

OpenStudy (anonymous):

so it would be 1/2*4 =2 ??

OpenStudy (andriod09):

kind of. just look.

OpenStudy (andriod09):

coin 1 \[\frac{1} {2}\] coin 2 \[\frac{1} {2}\] coin 3 \[\frac{1} {2}\] coin 4 \[\frac{1} {2}\]

OpenStudy (anonymous):

wouldn't be 50% percent ?

OpenStudy (andriod09):

that is the chance of getting heads on all for of them. so then it would be: \[\frac{4} {2}\] so yea, you were right, my bad. cuz' when you simplify it would become: \[\frac{2} {1}\]

hartnn (hartnn):

when 4 coins are tossed only one event does not have head which is TTTT. T=tails so the probability will be 15/16

OpenStudy (anonymous):

no its 15/16 .. binomial .. u get 1 head by this no. of wasys = 4C1\[\left( \frac{ 1 }{ 2 } \right)^{4}\] u get 2 head by this no. of wasys = 4C2\[\left( \frac{ 1 }{ 2 } \right)^{4}\] u get 3 head by this no. of wasys = 4C3\[\left( \frac{ 1 }{ 2 } \right)^{4}\] u get 4 head by this no. of wasys = 4C4\[\left( \frac{ 1 }{ 2 } \right)^{4}\] total no. of ways = (adding the above quantities) = 15/16

OpenStudy (anonymous):

its 4C1 * (1/2)^4

OpenStudy (andriod09):

okay, im in algebra 3, i just finished probablility, how dahell did you get \[\frac{15}{16}\]

OpenStudy (anonymous):

i m in 10+2 ... man this is the basic problem ... probaly the first question of the topic..

OpenStudy (anonymous):

google binomial distribution and then u can study it .. may be helpful..